What volume of 0.089 M HCl must be added to 765 mL of 0.123 M Na2CO3 to obtain a 9.7 pH?
Answer #1

The answer is given in the attachment.Answer Let us take, vme volume of 0.089(M) te must be added. Nag CO ₂ + Heu Nace & H₂CO3 No. of millimoles of Nao co = 765* 00.089v I 194.095 -0.089 = -0.62 -0.62 - 0.089 v = (94.095 -0.089V) x 10 3 0.089 + 0.02135 v = 2.01 - 0.11035 v = 2.01 v 18.2